If the probability of a product having no defects is 0.925. What is the probability that a product will have at

   least one defect?
Answer #1

The probability of a product having no defects and the probability that a product will have at least one defect are complementary. Hence, the probability that a product will have at least one defect = 1 - P(no defects) = 1 - 0.925 = 0.075